# Characterizations of amenability for noncommutative dynamical systems and Fell bundles

### v1

@@ 5.5. Weak containment, nuclearity and amenability
The first claim is [17, Corollary 5.6]. Since $(\mathcal{B}_{N})_{H}=\mathcal{B}_{H}$ and we have conditional expectations, $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{N})$ and $E_{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}(\mathcal{B}_{N})$, we may assume that $N=G$ (i.e. $H$ is normal in $G$). We go a step further and use a trick of [16] to reduce the proof to $H=\{e\}$.
@@
Take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B})\to\mathbb{B}(X)$ with faithful integrated form. Via the isomorphism $L^{1}(\mathcal{L})\cong L^{1}(\mathcal{B})$, $T$ induces a *-representation $T^{1}\colon\mathcal{L}\to\mathbb{B}(X)$, see [16, VIII 15.9], and $T^{1}$ determines a unique *-representation $\dot{T}\colon{\mathcal{B}}_{/H}\to\mathbb{B}(X)$ such that $\dot{T}\circ\rho=T^{1}$. Moreover, the integrated forms of $\dot{T}$ and $T$ are the same up to the isomorphism of (5.4). Since the integrated form of $\dot{T}$ is faithful, the restrictions of $\dot{T}$ to each one of the fibers is isometric. Hence, for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$, $\|\rho(\Phi\circ f)\|=\|\dot{T}_{\Phi\circ f}\|$.
Take a nondegenerate dilation of $T\circ\Phi$, $S\colon\mathcal{B}\to\mathbb{B}(Y)$, and $V\in\mathbb{B}(X,Y)$ such that $T\circ\Phi(b)=V^{*}S_{b}V$ and $\|V\|\leq 1$. Using the explicit descriptions $T^{1}$ and $S^{1}$ of [16, VIII 15.9], a straightforward computation reveals that for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$, $T^{1}_{\Phi\circ f}=V^{*}S^{1}_{f}V=V^{*}\dot{S}_{\rho(f)}V$ and it follows that $\|\rho(\Phi\circ f)\|\leq\|\rho(f)\|$. Since $\rho(C_{c}(\mathcal{B}))$ is dense in $\dot{B}_{u}$, there exists a unique family of linear contractions $\{\dot{\Phi}_{u}\colon\dot{B}_{u}\to\dot{B}_{u}\}_{u\in G/H}$ whose union ${\Phi}_{/H}\colon{\mathcal{B}}_{/H}\to{\mathcal{B}}_{/H}$ satisfies ${\Phi}_{/H}\circ\rho(f)=\rho(\Phi\circ f)$, for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$. For all $f\in C_{c}(\mathcal{B})$ we have ${\Phi}_{/H}\circ(\rho\circ\tilde{f})(u)={\Phi}_{/H}\circ(\rho(f|_{u}))=\rho((\Phi\circ f)|_{u})=\widetilde{\Phi\circ f}(u)$. Considering the family $\Gamma:=\{\rho\circ\tilde{f}\colon f\in C_{c}(\mathcal{B})\}$ and using [16, II 13.16] we get that ${\Phi}_{/H}$ is continuous and it is ccp because $T\circ{\Phi}_{/H}(z)=V^{*}\dot{S}_{z}V$. Notice that ${\mathop{\rm supp}}({\Phi}_{/H})$ is contained in the projection of ${\mathop{\rm supp}}(\Phi)$ into $G/H$.
@@ Appendix A Tensor products of equivariant cp maps between $\mathrm{W}^{*}$-algebras
Assume that $1_{X}\in P$. We know that $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq\mathbb{B}(Z)\bar{\otimes}N$. Combining Lemma A.4 with [22, Corollary 5.9] we obtain $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq(P^{\prime}\bar{\otimes}\mathbb{C}1_{Y})^{\prime}=P\bar{\otimes}\mathbb{B}(Y)$. Now that we know
@@
we can use [22, Corollary 5.10] to get $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq P\bar{\otimes}N$.
@@
All we have done so far can be repeated flipping the tensors, we mean using $M\bar{\otimes}\mathbb{B}(Z)$ instead of $\mathbb{B}(Z)\bar{\otimes}M$, so we get a ccp map $\pi\bar{\otimes}\operatorname{id}\colon M\bar{\otimes}\mathbb{B}(Z)\to N\bar{\otimes}\mathbb{B}(Z)$. At this point we have all the necessary ingredients to prove equivariant versions of some results by Nagisa and Tomiyama [20, Theorem 3], [23, Theorem 4].

### v2

@@ 5.5. Weak containment, nuclearity and amenability
The first claim is [18, Corollary 5.6]. Since $(\mathcal{B}_{N})_{H}=\mathcal{B}_{H}$ and we have conditional expectations, $E\colon C^{*}(\mathcal{B})\to C^{*}(\mathcal{B}_{N})$ and $E_{\mathop{\rm r}}\colon C^{*}_{\mathop{\rm r}}(\mathcal{B})\to C^{*}_{\mathop{\rm r}}(\mathcal{B}_{N})$, we may assume that $N=G$ (i.e. $H$ is normal in $G$). We go a step further and use a trick of [17] to reduce the proof to $H=\{e\}$.
@@
Take a nondegenerate *-representation $T\colon C^{*}(\mathcal{B})\to\mathbb{B}(X)$ with faithful integrated form. Via the isomorphism $L^{1}(\mathcal{L})\cong L^{1}(\mathcal{B})$, $T$ induces a *-representation $T^{1}\colon\mathcal{L}\to\mathbb{B}(X)$, see [17, VIII 15.9], and $T^{1}$ determines a unique *-representation $\dot{T}\colon{\mathcal{B}}_{/H}\to\mathbb{B}(X)$ such that $\dot{T}\circ\rho=T^{1}$. Moreover, the integrated forms of $\dot{T}$ and $T$ are the same up to the isomorphism of (5.4). Since the integrated form of $\dot{T}$ is faithful, the restrictions of $\dot{T}$ to each one of the fibers is isometric. Hence, for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$, $\|\rho(\Phi\circ f)\|=\|\dot{T}_{\Phi\circ f}\|$.
Take a nondegenerate dilation of $T\circ\Phi$, $S\colon\mathcal{B}\to\mathbb{B}(Y)$, and $V\in\mathbb{B}(X,Y)$ such that $T\circ\Phi(b)=V^{*}S_{b}V$ and $\|V\|\leq 1$. Using the explicit descriptions $T^{1}$ and $S^{1}$ of [17, VIII 15.9], a straightforward computation reveals that for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$, $T^{1}_{\Phi\circ f}=V^{*}S^{1}_{f}V=V^{*}\dot{S}_{\rho(f)}V$ and it follows that $\|\rho(\Phi\circ f)\|\leq\|\rho(f)\|$. Since $\rho(C_{c}(\mathcal{B}))$ is dense in $\dot{B}_{u}$, there exists a unique family of linear contractions $\{\dot{\Phi}_{u}\colon\dot{B}_{u}\to\dot{B}_{u}\}_{u\in G/H}$ whose union ${\Phi}_{/H}\colon{\mathcal{B}}_{/H}\to{\mathcal{B}}_{/H}$ satisfies ${\Phi}_{/H}\circ\rho(f)=\rho(\Phi\circ f)$, for all $f\in\uplus_{u\in G/H}C_{c}(\mathcal{B}_{u})$. For all $f\in C_{c}(\mathcal{B})$ we have ${\Phi}_{/H}\circ(\rho\circ\tilde{f})(u)={\Phi}_{/H}\circ(\rho(f|_{u}))=\rho((\Phi\circ f)|_{u})=\widetilde{\Phi\circ f}(u)$. Considering the family $\Gamma:=\{\rho\circ\tilde{f}\colon f\in C_{c}(\mathcal{B})\}$ and using [17, II 13.16] we get that ${\Phi}_{/H}$ is continuous and it is ccp because $T\circ{\Phi}_{/H}(z)=V^{*}\dot{S}_{z}V$. Notice that ${\mathop{\rm supp}}({\Phi}_{/H})$ is contained in the projection of ${\mathop{\rm supp}}(\Phi)$ into $G/H$.
@@ Appendix A Tensor products of equivariant cp maps between $\mathrm{W}^{*}$-algebras
Assume that $1_{X}\in P$. We know that $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq\mathbb{B}(Z)\bar{\otimes}N$. Combining Lemma A.4 with [24, Corollary 5.9] we obtain $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq(P^{\prime}\bar{\otimes}\mathbb{C}1_{Y})^{\prime}=P\bar{\otimes}\mathbb{B}(Y)$. Now that we know
@@
we can use [24, Corollary 5.10] to get $(\operatorname{id}\bar{\otimes}\pi)(P\bar{\otimes}M)\subseteq P\bar{\otimes}N$.
@@
All we have done so far can be repeated flipping the tensors, we mean using $M\bar{\otimes}\mathbb{B}(Z)$ instead of $\mathbb{B}(Z)\bar{\otimes}M$, so we get a ccp map $\pi\bar{\otimes}\operatorname{id}\colon M\bar{\otimes}\mathbb{B}(Z)\to N\bar{\otimes}\mathbb{B}(Z)$. At this point we have all the necessary ingredients to prove equivariant versions of some results by Nagisa and Tomiyama [22, Theorem 3], [25, Theorem 4].